OpenStudy (anonymous):

find an equation of the line containing the given pair of points (-2,-2) and (-4,-9) y=

OpenStudy (anonymous):

y-y1=m(x-x1) i know you have to use that equation

OpenStudy (anonymous):

Maybe you can use (y-y1)/(y2-y1) = (x-x1)/(x2-x1) if that's easier

OpenStudy (anonymous):

First find the gradient (y1-y2)/(x1 -x2) = (-2--9)/(-2 --4) = 7/2 Therefore we know y = 7x/2 + c Put in a vlue of x and y from equation 1 as it is the easiest to work with -2 = 7(-2)/2 + c giving -2 = -7 +c adding 7 to both sides we get... 5 = c ...and now we can write the equation y -= 7x/2 + 5